OpenStudy (anonymous):

How to solve an exponential equation? 5^(x+6)=7

OpenStudy (anonymous):

I know I have to take the log of both sides.

OpenStudy (anonymous):

\[log_5(7) = (x+6)\]

OpenStudy (anonymous):

I'm confused. where did the 5 with the x+6 go?

OpenStudy (anonymous):

To use a calculator with common log of base-10, do log(7) divided by log(base). in this case \[\frac{log7}{log5}\]

OpenStudy (anonymous):

\[log_5(some number)\] is asking, what power of 5 is equal to somenumber

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so \[\frac{log7}{log5} = x+6\] and solve for x

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

could you help me with one more problem?

OpenStudy (anonymous):

Just be sure to remember that trick. Sometimes we have to change our log base, and a lot of calculators don't have an easily accessible button for this. TI calcs may have it buried in their menu systems but I never bother cause it takes too long.

OpenStudy (anonymous):

sure, if i am able to

OpenStudy (anonymous):

ok. \[\log_{8}(x^2-7x) = 1 \]

OpenStudy (anonymous):

This is saying 8 raised to the power 1 is equal to x^2-7x.

OpenStudy (anonymous):

so quadratic formula?

OpenStudy (anonymous):

\[log_{10}(100) = ?\]

OpenStudy (anonymous):

Not related, but solve it.

OpenStudy (anonymous):

?=2

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

So the little base number is always raised to the power on the right side of the equal, and the answer is inside the ( ).

OpenStudy (anonymous):

Just remember that.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[x^2-7x = 8^1\]

OpenStudy (anonymous):

\[x^2 - 7x - 8 = 0\]

OpenStudy (anonymous):

factor and solve for x, setting each term equal to zero

OpenStudy (anonymous):

\[x^2-7x-8 = (x-8)(x+1)\]

OpenStudy (anonymous):

set each term equal to zero and solve for x. To factor \[x^2-7x-8\] i found two numbers that multiply to get -8 and add to get -7

OpenStudy (anonymous):

you could use quadractic equation but that is slower

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[[(x-8)=0 ]\wedge [(x+1) = 0]\]

OpenStudy (anonymous):

x = 8 and x = -1

OpenStudy (anonymous):

\ok. that makes much more sense than when my teacher explained it

OpenStudy (anonymous):

\[\wedge\] is a symbol for and

OpenStudy (anonymous):

cool :) that's what OS is for

OpenStudy (anonymous):

wait till you get to calclulus, then you start enjoying math cause it's useful and not boring lol. but u gotta master algebra 1st so you're doing good

OpenStudy (anonymous):

haha. this is calculus, at least, that's what the class is called

OpenStudy (anonymous):

pre-calculus?

OpenStudy (anonymous):

yea, sorry. pre-calc

OpenStudy (anonymous):

almost there!

OpenStudy (anonymous):

not taking calculus next year at all.

OpenStudy (anonymous):

no, my teacher made it clear that if we don't get this then we should just work at McD's/ so thats what im doing.

OpenStudy (anonymous):

oh man you should that's where everything starts to get interesting

OpenStudy (anonymous):

it's true

OpenStudy (anonymous):

(minimum wage waste-of-life jobs without advanced analytical skills)

OpenStudy (anonymous):

alright message anytime. best wishes, good job :)

OpenStudy (anonymous):

thank u!